A town’s average snowfall is 49 inches per year with a standard deviation of 5 inches. The distribution is symmetric and bell shaped. According to Empirical Rule, what amount of snowfall would you expect to be unusual for this town?

A) 53 inches

B) 63 inches

C) 35 inches

D) none of the given answers
Answer #1

The correct options are B and C

Let's write the given information.

mean = \mu = 49

standard deviation = \sigma = 5

According to the empirical rule 95% data lies within the 2\sigma limits. That is in usual range.

Therefore limits of usual range are as follows:

Lower limit = \mu - 2*\sigma = 49 - 2*5 = 39

Upper limit = \mu + 2*\sigma = 49 + 2*5 = 59

Therefore, the values which are outside (39 , 59) this intervals are unusual values.

So that 63 and 35 are unusual values.